The time taken by Mars to revolve round the sun is 1.88 years. Find the ratio of average distance between Mars and the sun to that between the earth and the sun.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. According to Kepler’s laws of planetary motion, T^2 ∝ R^3 (R base ms/R base es)^3 = (1.88/1)^2 ∴ R base ms/R base es (1.88)^2/3 = 1.52